OpenStudy (anonymous):

How do you solve the problem Pleaseeeeee help The number of drive-in movie screens has declined steadily in the united states. Pennslyvania and Ohio had the most remaining drive-in movie screens in 2007, Pennsylvania had 2 more screens than Ohio, and there were 68 screens total in the two states. How many drive- in movie screens remained In each state?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

P: number of screen in Pennsylvania O: number of screen in Ohio O + P = 68 and we know that p = o+2 O + O + 2 =68 2O + 2 =68 2O = 66 O = 33 P = 33 +2 =35